OpenStudy (anonymous):

OMG BULLY ON OPEN STUDY HELPPP!!!!! sorry for the title I just need your attention... Which is a solution of the equation? 2x – y = 0 A. (1, 2) B. (3, –1) C. (2, –4) D. 3

OpenStudy (anonymous):

this is linear equations, and inequalites (spell check)

OpenStudy (anonymous):

medal go 2 to the right one

OpenStudy (anonymous):

i have more btw

OpenStudy (accessdenied):

Is it the kind of attention you want, though? ;[ Anyways, if you have a solution to an equation, it can be substituted into the equation to give some sort of obviously true equality: like 5 = 5, you can't really say 5 stopped being 5. Like for (1, 2) A, 2x - y = 0 We could put x=1 and y=2 as with the points. 2(1) - (2) = 0 Does this turn out to 0 = 0 when computed?

OpenStudy (anonymous):

@AccessDenied um what does "computed" mean

OpenStudy (accessdenied):

Same as calculate... like you multiply everything out 2(1) = 2 and then do the subtraction, the end result.

OpenStudy (anonymous):

ok is it 1,2??????????

OpenStudy (accessdenied):

Yes. This one is easier, the answer was put first. If it hadn't been, say B(3, -1) came before it... this is an example where it does not work: 2(-1) - (3) = 0 -2 - 3 = 0 -5 = 0 <-- Now you'd never say these entities were the same. So B is not the answer.